Scott has $500 in the bank. He withdraws the same amount of money each week for 12 weeks, at which point he has $20 left in the bank. How much money did he withdraw each week?

Answer by jim_thompson5910(35256)

500 - 12x = 20

-12x = 20 - 500

-12x = -480

x = -480/(-12)

x = 40

So he withdraws $40 each week.